OpenStudy (anonymous):

Just testing to see if I can use latex on this site. [tex]/frac{1}{2}[/tex]

OpenStudy (anonymous):

is that a no?

OpenStudy (anonymous):

[math]\frac{1}{2}[/math]

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

\[\frac12\]

OpenStudy (anonymous):

\frac{1}{2}

OpenStudy (anonymous):

what command did u use waylander?

OpenStudy (anonymous):

\frac12

OpenStudy (anonymous):

umm not that

OpenStudy (anonymous):

was playing

OpenStudy (anonymous):

[tex]\frac{1}{2}[/tex]

OpenStudy (anonymous):

there is a equation tab below to do that

OpenStudy (anonymous):

\[1/2\]

OpenStudy (anonymous):

can you use commands rather than the editor?

OpenStudy (anonymous):

yea just a \ before n after

OpenStudy (anonymous):

\frac{1}{2}\

OpenStudy (anonymous):

\ [ 1/2 \ ] post that without space..

OpenStudy (anonymous):

\[1/2\]

OpenStudy (anonymous):

cool, thanks.

OpenStudy (anonymous):

\[\frac{1}{2}\]

OpenStudy (anonymous):

\[\frac{(x-1)^2}{(x+2)^2}\]

OpenStudy (anonymous):

\frac12\

OpenStudy (anonymous):

\(1/2)\

OpenStudy (anonymous):

\[1/2\]

OpenStudy (anonymous):

\ [ \ frac{2}{3} \ ] without spaces

OpenStudy (anonymous):

\[\frac{2}{3}\]

OpenStudy (anonymous):

\[\frac23\]

OpenStudy (anonymous):

dont need the brackets around the numerator/denominator

OpenStudy (anonymous):

\[\frac23\ ]

OpenStudy (anonymous):

\[\frac{log_a b}{\sqrt{2}}\]

OpenStudy (anonymous):

ahh ok

OpenStudy (anonymous):

\[\frac{45}{50}\]

OpenStudy (anonymous):

you do if the numbers are two digit or you have more than one term